### v5

# Sensitive parameter dependence of the self-similar dynamics in the bubble breakup in a confined geometry

Abstract
Formation of a fluid drop, such as observed in dripping faucet, has been extensively studied for understanding singular dynamics widely observed in nature. The singular dynamics is often self-similar, i.e., the shape at different times are collapsed onto a master curve after rescaling. Here, we investigate the bubble breakup in an original setup, where a metal disk entrains air into viscous fluid with lubricating films on its surfaces in a confined cell, focusing on the dependence of the dynamics on the film thickness $e$. As a result, we find out the dynamics is self-similar as expected, but the shape of the master curve and the scaling exponents characterizing the self-similar dynamics are revealed to be sensitively dependent on $e$. We discuss analogy between the present study and critical phenomena in thermodynamic transitions, to suggest a promising future direction of the singular dynamics.
I Introduction
Formation of a fluid droplet we observe on a daily basis, e.g., in a dripping faucet, has been extensively studied as representatives for singular dynamics widely observed in nature [1, 2], from gravitational collapse of stars [3] to cell aggregation [4]. Near the singularity, the shapes after rescaling often become independent from time collapsing on a master curve, i.e., the dynamics becomes self-similar [5]. Examples of the self-similar dynamics found in the formation of a fluid drop include breakup of a liquid drop in air, where inertia, capillary, and viscosity compete among others [6, 7], breakup of a bubble when high viscosity of the surrounding fluid competes with capillarity [8], and breakup of a bubble in a tube [9]. In all these examples, the self-similar dynamics of the breakup of a fluid drop is axisymmetric.
Recently, however, an example of the self-similar drop breakup without axisymmetry was reported [10] in a confined geometry and the breakup was further revealed to be suppressed if we further confined the system [11]. These unexpected findings were observed in an original system, in which a metal disk of radius $R$ ($\simeq$ cm) and of thickness $D_{0}$ ($\simeq$ mm) drags air into a viscous liquid phase with lubricating film of thickness $e$ ($\simeq$ mm) on the disk surfaces in a cell of thickness $D_{0}+2e$. However, in these previous studies, the shape dynamics was investigated mainly at $e=0.5$ mm. Here, we investigate the dependence of the shape dynamics on the film thickness $e$ for a fixed $D_{0}$ to find out that the shape dynamics changes sensitively with $e$ and the scaling exponents characterizing the self-similar dynamics are systematically dependent on $e$. This is in contrast with the previous studies [10, 11] in that the corresponding scaling exponents were found to be always the same value even when $D_{0}$ and $R$ were changed. We stress analogy between the self-similar dynamics and the scaling hypothesis in critical phenomena in thermodynamic transitions [12, 13]. This analogy suggests a promising avenue for the future study of singular dynamics by exploring symmetry in confined geometry.
Figure 1: (a) Experimental setup. A metal disk of thickness $D_{0}$ ($=1$ mm) and radius $R$ (101010~{}10to $12.5$ mm) falls in the cell of thickness $D$ ($2$ to $6$ mm) filled with a viscous liquid of kinematic viscosity $\nu$ (100 to 1000 cS). The disk entrains air into the liquid, which finally detaches from the disk. The difference between $D$ and $D_{0}$ defines the liquid film thickness $e$. (b) Snapshots just before and at breakup illustrating the setting of axes, in the case with topology change for $(R,D_{0},e,\nu)=(10,3,0.5,100)$ in mm or cS.
Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to detachment of air from the disk for $R=10$ mm, $D_{0}=1$ mm, $\nu=100$ cS. The liquid film thickness $e$ are 1, 1.5, and 2 mm, respectively, in (a) to (c). The time label 0 ms corresponds to $t=t_{d}$ defined in the text. In the three rightmost front-view photos reveal that a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a): ”the critical breakup” lies between $e=$ $1$ and 1.5 mm. Side-view shots near the detachment (0 ms) are separated by 6 ms, but are not synchronized with the front-view shots (a set of side-view snapshots are obtained from an experiment performed on a day different from the day on which the corresponding front-view snapshots but conducted for the same parameters).
II Results
II.1 Experimental
In Fig. 1 (a), we show our experimental setup with explanation. The ranges of characteristic lengths, the radius and thickness of disk and the cell thickness, are as follows: $R=10-12.5$ mm, $D_{0}=1$ mm, $D=2-6$ mm. The cell width and height are much larger than the length scales $R,D_{0}$, and $D$ (typically 9 and 12 cm, respectively). We use polydimethylsiloxane (PDMS) for viscous liquid, where the range of kinematic viscosity $\nu=\eta/\rho$ is 100-1000 cS. The density $\rho$ and the surface tension $\gamma$ are slightly depending on viscosity $\eta$ ($\rho\simeq$ $0.97$ g/cm3 and $\gamma\simeq 20$ mN/m). The density $\rho_{s}$ of the metal disk (SUS430) is 7.7 g/cm3 with the density difference $\Delta\rho=\rho_{s}-\rho$. The cell is fabricated with acrylic plates of thickness 5 mm, using acrylic spacers whose thickness defines the cell thickness $D$.
To obtain reproducible results, we set a gate at the top of cell by gluing a pair acrylic plates of thickness very close to $e=(D-D_{0})/2$, one for the back surface of the front cell plate and the other for the front surface of the back cell plate, to make the gap at the gate close to the disk thickness $D_{0}$. This gate helps to make the thickness of two liquid films between the surfaces of the disk and cell precisely equal to $e$ (see the Side View in Fig. 1 (a)). We fall the disk so that the initial speed of the disk is zero, i.e., the bottom of the disk is in contact with the interface at zero velocity at the entry. The disk surface is coated with a very thin layer of the same liquid as the one in the cell, by once dipping the disk into the liquid and then removing the liquid well with liquid-absorbing paper, to guarantee the zero static contact angle. We record the shape change of the air-liquid interface with a high-speed camera (FASTCAM Mini UX 100, Photron) with a lens (Micro NIKKOR 60 mm f2.8G ED, Nikon). The range of frame per second (fps) is 1000-2000. The images are analyzed with Image J and self-made Python codes.
II.2 Shape of the air-liquid interface $h(z,t)$
In Fig. 1 (b), we explain the setting of coordinates in the present study (we set $z_{c}=0$, so that $z-z_{c}=z$ and $z_{m}-z_{c}=z_{m}$). The air-liquid interface formed by air entrained by the disk is seen as the dark curve with a finite width. The width reflects the three-dimensional characters of the interface: the interface is concave or convex. The region where the width is thin, the interface is almost flat. We track the right and left inner edges of the thick curve and describe them by $x=h(z,t)$ and $-h(z,t)$, respectively, except for the case of $e=0.5$ mm, for which we track instead the outer edges. We determine whether we track the inner or outer edge such that we can have a better collapse of shapes at different time after rescaling, which we discuss below. We call the minimum of the function $h(z,t)$ with respect to $z$ ”the constriction point” at which $(x,z)=(h_{m}(t),z_{m}(t))$, i.e., $h_{m}(t)=h(z_{m}(t),t)$.
II.3 Dynamics sensitive to the film thickness $e$
In Fig. 2, we show snapshots of air-detachment transitions near the air-solid contact for $e=1.0$, 1.5, and 2 mm, with the other parameters, $R$, $D_{0}$, and $\nu$ fixed (we have tried to analyze the case of $e=2.5$ mm, only to fail because the inner edge was not clear). In three rightmost front-view photos (labeled as 6 ms), detachment without topological change is observed for $e=1$ mm but that with topological change (i.e., breakup) for $e=1.5$ and $2$ mm: a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a). In other words, ”the critical breakup” lies between $e=$ $1$ and 1.5 mm.
The origin ”0 ms” of the time label ($t-t_{d}$) given in ms in Fig. 2 and figures below for the discussion of self-similar dynamics is set for the snapshots either just before or after the detachment and thus can be slightly deviated from the real detaching time $t_{d}^{\prime}$ with an error, which should be at most 0.5 ms when images are recorded at 2000 fps.
In the previous studies [10] and [11], the cases of $D_{0}\gtrsim\kappa^{-1}$ and $D_{0}\lesssim\kappa^{-1}$ with $\kappa^{-1}=\sqrt{\gamma/(\rho g)}\simeq 1.8$ mm are distinguished by the terms sheet-forming and corn-forming detachments, and the former and latter cases are respectively regarded as the cases with and without topology transition and with and without appearance of the constriction point. However, a careful examination on the $e$-dependence in the present study reveals that the previous distinction is applicable only for limited cases: in the present case, we observed both the cases with and without topology transition for $D_{0}\lesssim\kappa^{-1}$ and detachments shown in Fig. 2 are difficult to categorize by the sheet-forming or corn-forming case because it looks like the corn-forming case but the constriction point does appear.
II.4 Dynamics of characteristic length scales
In Fig. 3 (a1) to (a4), we show the dynamics of three characteristic length scales, $2h_{m}(t)$, $z_{m}(t)$, and $z_{G}(t)$, for different film thickness $e$. The quantity $z_{G}(t)$ is the position of the center of gravity of the disk measured from that at the critical time $t=t_{c}$, which slightly differs from $t_{d}$ [$z_{G}(t_{c})=0$ by definition]. The critical time $t_{c}$ is determined by extrapolation as the time when $h_{m}$ becomes zero. The critical space-time point is represented as $(x,z,t)=(x_{c},z_{c},t_{c})$ with $x_{c}=h_{m}(t_{c})$ and $z_{c}=z_{m}(t_{c})$.
[…]
Plot (a1) shows that three length scales are identical, i.e.,
$2h_{m}(t)=z_{m}(t)=z_{G}(t),$ (1)
at $e=0.5$ mm as also observed in different parameter ranges [10, 11]. In the present case, however, as shown in plots (a2) to (a4), they grow into different length scales as $e$ increases.
We regroup and show plots in (a) on a log-log scale in (b), which suggest the existence of scaling laws for these length scales, although the range of scaling is rather limited to nearly over, or slightly less than, one order of magnitude. In (b1), since we expect $z_{G}$ linearly scales with $t_{c}-t$ as discussed below, we fit the data by a linear function $y=ax$ with a fitting parameter $a$ in the region indicated by a pair of cross marks. The fitting line with slope one on the log-log scale thus obtained are shown in (b1). In (b2) and (b3), we fit the data by $y=ax^{b}$ with two fitting parameters $a$ and $b$ in the region indicated by a pair of cross marks. The fitting line with slope $b$ thus obtained are shown in (b2) and (b3). The log-log plots in (b1) to (b3) suggest, in addition to the linear scaling for $z_{G}$, the following scaling laws for $h_{m}$ and $z_{m}$:
$h_{m}=c_{1}t^{\prime\beta}\text{ and }z_{m}=c_{2}t^{\prime\Delta}$ (2)
with a time label $t^{\prime}$ (which is positive at times before the critical time $t=t_{c}$):
$t^{\prime}=t_{c}-t.$ (3)
The exponents $\beta$ and $\Delta$ thus obtained are summarized in Tab. 1, which shows a systematic dependence of the exponents on $e$. This is in contrast with the previous studies [10, 11], where $\beta$ and $\Delta$ are always one even if $D_{0}$ and $R$ are changed.
Table 1: Exponent $\beta$ and $\Delta$ determined from Fig. 3 by fitting the data with the functions, $h_{m}(t)\sim(t_{c}-t)^{\beta}$ and $z_{m}(t)\sim(t_{c}-t)^{\Delta}$, respectively. See the text for the determination of the exponent $\delta$.
$\begin{array}[c]{|c||c||c||c||c|}\hline\cr e\text{ (mm)}&\beta&\Delta&\delta&\Delta/(\beta\delta)\\ \hline\cr\hline\cr 0.5&1.0\pm 0.0&0.97\pm 0.06&1.01\pm 0.01&0.97\pm 0.06\\ \hline\cr 1.0&1.0\pm 0.0&0.79\pm 0.02&0.64\pm 0.00&1.2\pm 0.03\\ \hline\cr 1.5&0.93\pm 0.00&0.78\pm 0.01&0.61\pm 0.00&1.4\pm 0.02\\ \hline\cr 2.0&0.90\pm 0.01&0.77\pm 0.03&0.59\pm 0.00&1.5\pm 0.06\\ \hline\cr\end{array}$
We carefully selected the range of fitting indicated by a pair of cross marks for $h_{m}$ and $z_{m}$ to avoid including the data with large error bars and to use only the data with small error bars, by noting that (i) the error coming from the time resolution tends to become large near $t=t_{c}$ and (ii) the error tends to be large in $z_{m}$ compared with $h_{m}$. These points are visualized by the error bars shown for the data at $e=1.0$ mm in (b2) and (b3): The error bars in the horizontal axis are based on the time resolution ±$\pm 0.5$ ms at 2000 fps; The error bars in the vertical axis are on the spacial resolution corresponding to 1 pixel of the images for $h_{m}$ but possibly corresponding to several pixels for $z_{m}$ because of the elongation explained below. In accordance with Remark (i) and (ii), we can confirm the following in (b2) and (b3): the error bars in the horizontal axis is visible near $t=t_{c}$, while the error bars in the vertical axis are not visible for $h_{m}$ but significant for small $z_{m}$. The reason for large errors in $z_{m}$ is as follows. The constriction region tends to elongate near the breakup and the thin thread of air thus formed detaches from the disk when $h_{m}$ is not zero (see Fig. 6 below). Because of this elongation of the constriction region the determination of $z_{m}(t)$ tends to become difficult near breakup, which is reflected in larger standard deviation for $\Delta$ (compared with $\beta$) in Tab. 1.
As for the fitting of $z_{G}$, we justify the above linear fitting by Fig. 3 (c), which supports the following scaling law for $z_{G}(t)$ in agreement with the result established in a similar parameter range [11]:
$z_{G}(t)=v_{G}t^{\prime}\text{ for }t<t_{c},$ (4)
with a characteristic velocity scale
$v_{G}=k\Delta\rho gD_{0}e/\eta$ (5)
with the gravitational acceleration $g$. Note that the agreement between the data and the dotted line shown in (c) is demonstrated with using the value, $k=0.254$, which is the value obtained in the previous study [11].
II.5 Self-similarity in the interface shape dynamics
In Fig. 4 (a1), we show the interfacial shape before rescaling at different times for $e=1$ mm. Interface shapes after rescaling are reasonably collapsed well on a master curve, especially near the constriction point, as seen in (a2). The collapse is significant for shapes before $t-t_{d}=-15$ ms as shown in (a3).
Figure 4: Shape change for $e=1$, 1.5, 2.0, and 0.5 mm for $(R,D_{0},\nu)=(10,1,100)$ in mm or cS. Since the right and left interfaces are the mirror image of the other, only the right interface after averaging is shown. The error bars in capturing the interface are less than the size of markers. The time development of interfacial shape before [after] rescaling is respectively shown in (a1) [(a2) and (a3)] for $e=1$ mm. We see a better collapse for times not too close to $t=t_{c}$ (before $-9$ ms) [see (a3)]. The corresponding plots for $e=$1.5, 2.0, and 0.5 mm are shown respectively in (b) to (d). The data labeled $R^{\ast}$ in (a3) are obtained for different parameter set $(R,D_{0},\nu)=(12.5,1,100)$, which collapse well on the master curve [the good collapse seems limited for the upper branch (of our focus), i.e., the shape above the constriction point].
Similar collapses are shown in (b) and (c) for $e=1.5$ and 2.0 mm, respectively. As $e$ increases the constriction region tends to elongate near breakup as announced, which can be confirmed by comparing (b1) with (c1) (see also Fig. 6 below). This elongation is related to the above-mentioned criticality of breakup: topology does change for $e=1.5$ and 2 mm with accompanied by breakup but it does not for $e=1.0$ mm. It seems that topological change (i.e., breakup) tends to elongate the constriction region near $t=t_{d}$. Because of this elongation, duration of a good collapse tends to be limited as $e$ increases [see (a2) to (c3) and (a3) to (c3)].
Reproducibility and universality are also demonstrated in Fig. 4. The data labeled $R^{\ast}$ in (a3) imply that the corresponding data are obtained for different $R$ with the other parameters fixed. In Fig. 4 (a3), even these data obtained for different $R$ collapse well on the master curve, which demonstrates the master curve is independent from (i.e., universal for a change in) $R$ for a fixed $e$, indirectly representing a good reproducibility of the experiment.
In Fig. 4 (d1), we show the interfacial shape for $e=0.5$ mm, which is fundamentally different from cases of $e$ larger than 0.5 mm. In this case, as mentioned above, we need to track the outer edge instead of inner edge of the dark interface to observe collapse of the renormalized shape as seen in (d2) [and more clearly in (d3)]. In addition, the constriction point does not appear [the point $(h_{m},$ $z_{m})$ is defined as the right contact point (at $y=0$) of the interface with the disk] but the tip forms a corn. However, the corn may not be axisymmetric but elliptic as reported in our previous study [11].
The self-similarity demonstrated in Fig. 4 can be represented as
$h(t,z)=h_{m}(t)\Gamma(z/z_{m}(t))\equiv h_{m}(t)\Gamma(\xi),$ (6)
where the master curve is represented by the function $\Gamma(\xi)$ with $\xi=z/z_{m}(t)$. Figure 4 further suggests that the master curve is dependent on $e$. In fact, comparing the master curve in (a3) to (c3) with (d3), the shape difference is evident. Even among (a3) to (c3), we can regard the shape difference: as $e$ increases, the shape tends to be compressed in the vertical direction. In Fig. 4, we also notice that the master curve seems to be well described by
$\Gamma(\xi)=1+c(\xi-1)^{1/\delta},$ (7)
which scales as $\Gamma(\xi)\sim\xi^{1/\delta}$ for $\xi\gg 1$. This form and the exponent $\delta$ will be quantified below.
II.6 Identical shapes translating in time
We show in Fig. 5 (a1) to (d1) that the shapes are exactly identical for a given parameter set in the period not too close to $t=t_{c}$. They are just translating in space: when $z-z_{m}$ is plotted as a function of $h-h_{m}$, all the curves collapse onto a master curve. These master curves can be reasonably well described by the relation,
$h-h_{m}=c_{0}(z-z_{m})^{1/\delta}\equiv c_{0}\widetilde{z}^{1/\delta},$ (8)
as shown in Fig. 5 (a2) to (d2). The exponent $\delta$ is obtained by fitting the data in a range of small error bars shown by a pair of black cross marks in the plots (a2) to (d2), where the fitting lines thus obtained are shown. Numerical values of $\delta$ thus obtained are summarized in Tab. 1, which shows $\delta$ is systematically dependent on $e$ (In the previous study [11], the data corresponding at $e=0.5$ mm are fit by a curve with $\delta=0.75$, which is consistent with the present data: if we fit the data in the region between the pair of blue cross marks, we obtain $\delta=0.75$, which shows the present fitting focus on a region of small errors).
Figure 5: Translated shape functions for $e=1.0,1.5,2.0$, and 0.5 mm on linear scales in the period of good collapse [(a1) to (d1)] and on log-log scales for $-21$ to $-30$ ms with a fitting line obtained by fitting the data in the region between the two black crosses for $-21$ ms [(a2) to (d2)]. For all cases, $(R,D_{0},\nu)=(10,1,100)$ in mm or cS.
The exponent $\delta$ in Eq. (8) can be identified with $\delta$ in Eq. (7) because the former equation can be cast into the following form: $h/h_{m}=1+c^{\prime}(z_{m}^{1/\delta}/h_{m})(\xi-1)^{1/\delta}$ with $\xi=z/z_{m}$, where the collapse of the function at different times observed in Fig. 4 means that the coefficient $z_{m}^{1/\delta}/h_{m}$ is a constant,
$h_{m}\sim z_{m}^{1/\delta}.$ (9)
In other words, Eq. (8) reduces to Eq. (7) and thus $\delta$ in the both equations can be identified with each other.
Equation (9), together with Eq. (6), leads to the following relation:
$\Delta=\beta\delta.$ (10)
As seen in the column $\Delta/(\beta\delta)$ in Tab. 1, Eq. (10) is reasonably well satisfied at $e=1.0$ mm, and less satisfied as $e$ increases. Correspondingly, the spacial region of good collapse becomes more limited as $e$ increases.
II.7 Closed equation describing the shape
The translating interface observed in Fig. 5 can be described by the expression, $x=h_{m}+f(z-z_{m})$, from which we obtain $\partial h/\partial t=\dot{h}_{m}-f^{\prime}\dot{z}_{m}$ and $\partial h/\partial z=f^{\prime}$. These relations lead
$\frac{\partial h}{\partial t}+\dot{z}_{m}\frac{\partial h}{\partial z}=\dot{h}_{m}\text{.}$ (11)
This equation with Eq. (2) is a closed equation for the shape dynamics.
This equation can be derived from another expression of the translating interface: the velocity of points on the surface
$(u_{s}(t,z),v_{s}(t,z))\equiv(u(t,x=h(t,z),z),v(t,x=h(t,z),z))$ (12)
is independent of spacial coordinates. From this, we obtain the spacially constant velocity as the speed of the constriction point:
$(u_{s},v_{s})=(\dot{h}_{m},\dot{z}_{m}),$ (13)
for which Eq. (2) holds. Equation (11) can be derived as announced, if Eq. (13) is combined with the following equation of motion for the interface $h$ under Eq. (13):
$\frac{\partial h}{\partial t}+v_{s}\frac{\partial h}{\partial z}=u_{s}.$ (14)
This can be derived from the fact that a point on the interface $(x,z)$ with $x=h(t,z)$ translates with a velocity $(u_{s}(t,z),v_{s}(t,z))$ defined in Eq. (12), i.e., the point moves to the point $(x^{\prime},z^{\prime})=(x+u_{s}dt,z+v_{s}dt)$ with $x^{\prime}=h(t+dt,z^{\prime})$ after $dt$.
III Discussion
III.1 Universality and separation of scales
We show in Fig. 4 that the shape dynamics in the present parameter range is self-similar in limited ranges of time and space by demonstrating the collapse onto a master curve $\Gamma(\xi)$ by rescaling specified in Eq. (6), where $\Gamma(\xi)$ is dependent on $e$. From the viewpoint of separation of scales, we expect the present $\Gamma(\xi)$ is independent from $R$, as supported by the data labeled by $R^{\ast}$ in Fig. 4 (a3). This is because among the independent characteristic lengths $e$, $D_{0}$, and $R$, $R$ ($\sim 1$ cm) is the largest scale and thus is well separated from the characteristic scales for the self-similar dynamics. We may also expect $\Gamma(\xi)$ is independent from $\eta$ because the master curve $\Gamma(\xi)$ in the previous studies [10, 11] is always independent from $\eta$. In fact, we confirmed the master curve obtained for $(e,R,D_{0},\nu)=(0.5,10,1.5,1000)$ well collapses on the master curve in Fig. 4 (b3) obtained for $(e,R,D_{0},\nu)=(0.5,10,1.0,100)$ near the constriction point (in the region around $z/z_{m}=0$ to 1.4), which suggests the independence of $\Gamma(\xi)$ from $\eta$ and $D_{0}$. The independence from $D_{0}$ of $\Gamma(\xi)$ suggested here is also natural from the viewpoint of separation of scales since $D_{0}$ is the second largest scale. The independence from $R$ and $\eta$ of the $e$-dependent $\Gamma(\xi)$ suggested here is in accord with the dimensional analysis given below.
III.2 Comparison with the bubble breakup in non-confined geometry
The translation of the interface in time with keeping the same shape observed in the present study is a property shared by the breakup of a bubble in non-confined geometry as explained below. Under no confinement, the breakup occurs with the axisymmetry. Accordingly, the shape can be described in the cylindrical coordinate $(r,z,\theta)$ by $r=h(t,z)$, where the shape function $h(t,z)$ is independent of the $\theta$ coordinate from the axisymmetry. When a high viscosity $\eta$ of the surrounding fluid competes with capillary characterized by the surface tension $\gamma$, the dynamics is known to be described by the following equation (e.g., see Sec. 9.3.1 of [2]):
$\frac{\partial h(t,z)}{\partial t}=-\gamma/\eta$ (15)
Since the equation of motion of the interface in Eq. (14) is valid also in this case, Eq. (15) means the surface velocity is given as
$(u_{s},v_{s})=(-\gamma/\eta,0).$ (16)
This means that the shape is just translating in time in the horizontal direction as in the present case but with no translation in the vertical direction.
III.3 Dimensional analysis
We can physically understand how the scaling structure in Eq. (6) with Eqs. (1) and (4) at $e=0.5$ mm emerges in a natural manner. The key observation is the present problem can be regarded as finding a solution for Navier-Stokes equation for a viscous liquid, neglecting inertia and the role of air. To solve the problem we specify the boundary conditions, which are characterized by length scales $R$, $e$, and $D_{0}$. In addition, we can give a condition at $t=t_{c}$ to simplify the problem: the tip is moving at the velocity of the falling disk $v_{G}$. This velocity given in Eq. (5) is determined by a viscosity-gravity balance for a falling disk [11]: the viscous dissipation per time $\eta(v_{G}/e)^{2}R^{2}e$ occurring in the film of thickness $e$ whose volume scales as $R^{2}e$ balances with the gravitational energy change per time $\Delta\rho gR^{2}D_{0}v_{G}$. Furthermore, we note that the scaling structure in Eq. (6) with Eqs. (1) and (4) is characterized by dimensional parameters $\eta$, $\Delta\rho$ and $g$ alone, except for length scales $R$, $e$, and $D_{0}$, but is not involved with $\gamma$ and $\rho$, which again supports the dynamics governed by a viscosity-gravity balance. Accordingly, dimensionally, we expect
$h(t,z)=f(t^{\prime},z,\Delta\rho,\eta,g,R,e,D_{0})$ (17)
we have 9 dimensional variable, of which only 6 are independent, since the dimension of the unit of all the 9 quantities can be derived from the three fundamental units, kg, m, and s. From the Buckingham $\pi$ theorem [14], we expect a relation $\pi_{0}=\Xi(\pi_{1},\pi_{2},\ldots,\pi_{5})$, where $\pi_{i}$’s are 6 independent dimensionless variables and $\Xi$ is a dimensionless function. We select these dimensionless variables by focusing on 7 independent length scales $h$, $z$, $e$, $D_{0}$, $R$, $l=\eta/(\Delta\rho gt^{\prime})$ and $h_{m}\simeq z_{m}\simeq v_{G}t^{\prime}$, and set $\pi_{0}=h/h_{m}$ and $\pi_{1}=z/z_{m}$. To select the remaining 4 independent variables, we normalize the remaining 4 scales $e$, $D_{0}$, $R$, and $l$ by $z_{m}$:
$h=h_{m}\Xi(z/z_{m},e/z_{m},D_{0}/z_{m},R/z_{m},l/z_{m})$ (18)
Here, we may expect that near the breakup point where $h_{m}\simeq z_{m}$ is small so that $e/z_{m}$, $D_{0}/z_{m}$, $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ eventually approach towards infinity. This implies that if the function $\Xi$ remains finite and non-zero, it becomes independent of $e/z_{m}$, $D_{0}/z_{m}$, $R/z_{m}$ and $l/z_{m}$ to recover Eq. (6) with Eqs. (1) and (4) at $e=0.5$ mm.
If we assume Eq. (18) is valid even for $e=1.0$ to 2.0 mm, we understood the remaining dependence of the master curve $\Gamma$ on a small length scale $e$ (and possibly on $D_{0}$, which is comparable to $e$) from separation of scales. In the time domain in which the collapse is observed in Fig. 4, while $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ are large enough, $e/z_{m}$ is not large enough so that the $e$-dependence remains in $\Gamma$. The present dimensional argument thus suggests the independence from $R$ and $\eta$ of the $e$-dependent $\Gamma(\xi)$ as announced above. However, to justify Eq. (18) even for $e=1.0$ to 2.0 mm, we need to know the parameter dependence of the coefficients $c_{1}$ and $c_{2}$ in Eq. (2), which we will explore in a future study.
III.4 Analogy with critical phenomena
In critical phenomena in thermodynamic transitions, various thermodynamic quantities exhibit power laws near the critical point. Typical example is the ferromagnetic transition: the macroscopic magnetization $M(T,H)$ at zero magnetic field $H=0$ observed at low temperatures $T$ disappears at a critical temperature $T=T_{c}$, exhibiting a power law: $M(T,H=0)\simeq\Delta T^{\beta}$ with $\Delta T=\left|T-T_{c}\right|$. Here and hereafter, for simplicity, $T$ and $H$ are dimensionless temperature and magnetic field, respectively. The magnetization is an order parameter of critical phenomena, which is non-zero below $T_{c}$ and is zero above $T_{c}$ at $H=0$.
The power law is observed for different quantities. At $H=0$, the specific heat $C$ and the susceptibility $\chi$ behave as $C\simeq\Delta T^{-\alpha}$ and $\chi\simeq\Delta T^{-\gamma}$, while the magnetization $M$ behaves as $M\simeq H^{1/\delta}$ at $T=T_{c}$. The exponents such as $\alpha$ and $\beta$ are called the critical exponents and their values are governed by dimensionality $d$ and symmetry characterized by the number of components $n$ of the vector representing the order parameter. In other words, the exponents do not depend on some details such as the strength of the microscopic interaction and the structure of lattice. As a result, various materials characterized by the same $n$ and $d$ share the same critical exponents, forming a universality class. Namely, the same set of critical exponents universally describes critical behaviors such as $C\simeq\Delta T^{-\alpha}$ and $\chi\simeq\Delta T^{-\gamma}$ of different materials, if they belong to the same universality class although some quantities such as the critical temperature $T_{c}$ are material-dependent. Furthermore, the critical exponents satisfy scaling relations such as $\alpha+2\beta+\gamma=2$ and such relations are universally satisfied by the critical exponents of any universality classes. Such remarkable features can be understood based on the scaling hypothesis as explained in Appendix, and the scaling hypothesis naturally emerges in the renormalization group theory [12, 13].
The scaling hypothesis in the magnetic case can be expressed as
$M(T,H)\simeq\Delta T^{\beta}\Psi(H/\Delta T^{\Delta})\text{,}$ (19)
which bears a strong resemblance with the self-similar structure in Eq. (6) with Eq. (2). Time $t$ ($=t-t_{c}$) and position $z$ ($=z-z_{c}$) correspond to temperature $T$ and external field $H$, respectively, for ”the order parameter $h(t,z)$,” a quantity which is non-zero before $t_{c}$ but is zero after $t_{c}$ at $z=0$: the pre-breakup dynamics of our focus corresponds to the temperature range below $T_{c}$. The master curve $\Gamma(\xi)$ corresponds to the function $\Psi(\xi)$, which is called the scaling function and defines the critical exponent $\delta$ through $\Psi(\xi)\simeq\xi^{1/\delta}$ for $\xi\gg 1$ (and thus $\delta$ in Eq. (7) corresponds to the critical exponent $\delta$). The scaling exponents $\beta$ and $\Delta$ in Eq. (2) correspond to $\beta$ and $\Delta$ in Eq. (19), where the exponents $\beta$, $\Delta$, and $\delta$ in critical phenomena satisfy Eq. (10) as shown in Appendix.
In other words, the critical exponents $\beta$, $\Delta$, and $\delta$ in the present case are the ones summarized in Tab. 1. This table suggests that these three exponents are independent, where Eq. (10) is more violated as $e$ increases, and that that the exponents are monotonic functions of the thickness $e$. In addition, we may expect from the viewpoint of separation of scales and the dimensional analysis discussed above the exponents are universally shared by the systems with different $R$ and $\eta$ for a given $e$ (at $D_{0}=1$ mm). Namely, the present universality class may be determined by a continuous parameter $e$. This is in contrast with typical critical phenomena, in which universality class is determined by discrete parameters such as $n$ and $d$, although examples of critical exponents depending on continuous parameters are known when marginal operators play a role, which includes non-linear diffusion equations [15]. In the present case, a renormalization group analysis suggests a different scenario, which will be discussed elsewhere.
IV Conclusion
We investigate the pre-breakup dynamics of a bubble surrounded by a viscous liquid in a confined geometry focusing on the dependence on the lubricating film thickness $e$ and find that the dynamics is self-similar and that the scaling exponents characterizing the self-similarity is sensitively dependent on $e$. We elucidate similarity between the present singular dynamics and critical phenomena, showing that the counterpart of the critical exponents are dependent on a continuous variable $e$ and could be shared by a universality class formed by the systems with different $R$ and $\eta$ for a fixed $e$.
The analogy between the singular dynamics and critical phenomena elucidated in the present study underscores the importance of exploring symmetry in the future study of the singular dynamics by exploiting confined geometries for the following reasons. (1) In critical phenomena, which has propelled the development of the modern physics from soft and hard condensed matter to non-equilibrium systems and active matter [16, 17, 18, 19], dimensionality and symmetry have played a vital role in its development. (2) The singular dynamics is widely observed in nature and a variety of confined geometries are increasingly important in industrial applications and natural phenomena, from microfluidics [20, 21], petroleum industry [22], to geology [23]. In fact, the analogy with critical phenomena can be deepened further by developing a renormalization group analysis for the present singular dynamics, which will be discussed elsewhere.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596.
Appendix A Scaling relations in critical phenomena
Figure 6: Snapshots obtained at 2000 fps near the detachment point for $e=0.5$ mm (top) and $e=1.0$ mm (bottom) at $t-t_{d}=-1.5$ to $+1.5$ ms, respectively. The snapshots at the center correspond to $t-t_{d}=-0$ ms, where ”the order parameter” $h_{m}$ is not precisely zero.
We summarize here the basic knowledge on critical phenomena, focusing on the temperature range below $T_{c}$ ($T<T_{c}$). Excluding the exponent describing the spacial correlation, there are 4 exponents, $\alpha,\beta,\gamma$, and $\delta$, which are defined in the text as (c1) $C\simeq\Delta T^{-\alpha}$, (c2) $M\simeq\Delta T^{\beta}$, and (c3) $\chi\simeq\Delta T^{-\gamma}$ at $H=0$, together with (c4) $M\simeq H^{1/\delta}$ at $\Delta T=0$. However, among the four exponents, only two are independent. This results from the following thermodynamic relations, which are proven to be true by virtue of the renormalization group: The free energy $F(T,H)$ possesses the scaling structure (t1) $F(T,H)\simeq\Delta T^{2-\alpha}\Phi(H/\Delta T^{\Delta})$ near the critical temperature with a scaling function $\Phi(x)$, where thermodynamic quantities are derived from $F(T,H)$ as (t2) $C\simeq\partial^{2}F/\partial T^{2}$, (t3) $M\simeq\partial F/\partial H$, and (t4) $\chi\simeq\partial^{2}F/\partial H^{2}$. With these thermodynamic relations (t1) to (t4), the definitions of the critical exponents (c1) to (c4) result in the relations (a) $\Delta T^{-\alpha}\Phi(H/\Delta T^{\Delta})\simeq\Delta T^{-\alpha}$, (b) $\Delta T^{2-\alpha-\Delta}\Phi^{\prime}(H/\Delta T^{\Delta})\simeq\Delta T^{\beta}$ (i.e., another scaling function $\Psi(x)$ introduced above in Eq. (19) is expressed as $\Psi=\Phi^{\prime}$), (c) $\Delta T^{2-\alpha-2\Delta}\Phi^{\prime\prime}(H/\Delta T^{\Delta})\simeq\Delta T^{-\gamma}$ at $H=0$ and (d) $\Delta T^{2-\alpha-\Delta}\Phi^{\prime}(H/\Delta T^{\Delta})\simeq H^{1/\delta}$ at $\Delta T=0$. By requiring, based on consistency, that $\Phi(0)$, $\Phi^{\prime}(0)$, and $\Phi^{\prime\prime}(0)$ are non-zero finite value, while relation (a) simply means consistency, relations (b) and (c) lead to the following two independent relations, (A) $2-\alpha-\Delta=\beta$ and (B) $\beta-\Delta=-\gamma$, which result in the scaling relation $\alpha+2\beta+\gamma=2$ discussed in the text. Furthermore, relation (d) concludes, again based on consistency, the asymptotic behavior $\Psi(x)\simeq x^{1/\delta}$ for large $x$ with Eq. (10), if Eq. (19) is valid even for $x=H/\Delta T^{\Delta}$ is large. In this way, we can prove three independent relations (A), (B) and Eq. (10) for 5 exponents, $\alpha,\beta$, $\gamma$, $\delta$, and $\Delta$, which proves that there are only 2 independent exponents. However, Eq. (6), which corresponds to Eq. (19) in the magnetic case, holds only for relatively small $\xi=z/z_{m}(t)$, as seen in Fig. 4. This suggests breaking of Eq. (10), which allows the existence of three independent exponents, as implied in Tab. 1.
Appendix B Dynamics close to the breakup time
We compare in Fig. 6 the dynamics close to $t=t_{c}$, which shows qualitative differences for $e=0.5$ and 1.0 mm. In the former, sharp tip is formed, which detaches from the disk. In the latter, the constriction region elongates near $t=t_{c}$, forming a thin thread of air, which detaches from the disk when $h_{m}$ is not zero.

### v6

@@
-Sensitive parameter dependence of the self-similar dynamics in the bubble breakup in a confined geometry
+Breakup to non-breakup transition of air entrained into viscous liquid by a disk: analogy of the self-similar dynamics with critical phenomena
 Abstract
-Formation of a fluid drop, such as observed in dripping faucet, has been extensively studied for understanding singular dynamics widely observed in nature.
-The singular dynamics is often self-similar, i.e., the shape at different times are collapsed onto a master curve after rescaling.
-Here, we investigate the bubble breakup in an original setup, where a metal disk entrains air into viscous fluid with lubricating films on its surfaces in a confined cell, focusing on the dependence of the dynamics on the film thickness $e$.
-As a result, we find out the dynamics is self-similar as expected, but the shape of the master curve and the scaling exponents characterizing the self-similar dynamics are revealed to be sensitively dependent on $e$.
-We discuss analogy between the present study and critical phenomena in thermodynamic transitions, to suggest a promising future direction of the singular dynamics.
+Self-similarity in partial differential equations has been widely exploited to study many phenomena in physical sciences.
+We have studied the interfacial dynamics when air is entrained into viscous liquid by a disk in a confined geometry.
+In a previous study using an original experimental system, we found the sheet- and corn-forming regimes, in which a sheet and cone of air are respectively formed before air detaches from the disk.
+The sheet eventually breaks up but the corn, which appears when a bit more confined, does not.
+Here, we find a third regime, in which a corn eventually breaks up, by investigating different ranges of confining parameters: the transition from breakup to non-breakup can occur within the corn regime.
+Furthermore, with the data obtained in the third regime we deeply explore analogy with critical phenomena to find out that the counterpart of the critical exponents dependent on a length scale.
+Since the scale is a number not discrete but continuous, the present hydrodynamic analog suggests the existence of an uncountably infinite number of universality classes.
+The rich physics revealed in our study suggests a promising direction of the study of the self-similar dynamics: exploring analogy with critical phenomena, focusing on confined geometries in many natural and industrial phenomena.
 I Introduction
-Formation of a fluid droplet we observe on a daily basis, e.g., in a dripping faucet, has been extensively studied as representatives for singular dynamics widely observed in nature [1, 2], from gravitational collapse of stars [3] to cell aggregation [4].
-Near the singularity, the shapes after rescaling often become independent from time collapsing on a master curve, i.e., the dynamics becomes self-similar [5].
-Examples of the self-similar dynamics found in the formation of a fluid drop include breakup of a liquid drop in air, where inertia, capillary, and viscosity compete among others [6, 7], breakup of a bubble when high viscosity of the surrounding fluid competes with capillarity [8], and breakup of a bubble in a tube [9].
+When shapes as a function of $x$ defined by a solution $h(t,x)$ of a partial differential equations (PDE) at different $t$ collapse onto a single master curve, the dynamics is called self-similar.
+The self-similarity in PDEs has played a vital role in understanding many phenomena in physical sciences [1, 2], in particular, phenomena categorized as the singular dynamics [3, 4], which include fluid-jet formation [5, 6], drop coalescence [7, 8, 9], gravitational collapse of stars [10] and cell aggregation [11].
+Among them, the formation of a fluid droplet occurring, for example, in a dripping faucet, has been extensively studied.
+For the breakup of a liquid drop in air, a dynamical regime, in which inertia, capillary, and viscosity compete among others, is known [12, 13], while, for the breakup of a bubble, a regime, in which high viscosity of the surrounding fluid competes with capillarity, is found [14].
+The latter regime is further found to exhibit a crossover to another viscous-capillary regime in the breakup of a bubble in a tube [15].
 In all these examples, the self-similar dynamics of the breakup of a fluid drop is axisymmetric.
-Recently, however, an example of the self-similar drop breakup without axisymmetry was reported [10] in a confined geometry and the breakup was further revealed to be suppressed if we further confined the system [11].
-These unexpected findings were observed in an original system, in which a metal disk of radius $R$ ($\simeq$ cm) and of thickness $D_{0}$ ($\simeq$ mm) drags air into a viscous liquid phase with lubricating film of thickness $e$ ($\simeq$ mm) on the disk surfaces in a cell of thickness $D_{0}+2e$.
-However, in these previous studies, the shape dynamics was investigated mainly at $e=0.5$ mm.
-Here, we investigate the dependence of the shape dynamics on the film thickness $e$ for a fixed $D_{0}$ to find out that the shape dynamics changes sensitively with $e$ and the scaling exponents characterizing the self-similar dynamics are systematically dependent on $e$.
-This is in contrast with the previous studies [10, 11] in that the corresponding scaling exponents were found to be always the same value even when $D_{0}$ and $R$ were changed.
-We stress analogy between the self-similar dynamics and the scaling hypothesis in critical phenomena in thermodynamic transitions [12, 13].
-This analogy suggests a promising avenue for the future study of singular dynamics by exploring symmetry in confined geometry.
+Recently, however, an example of the self-similar drop breakup without axisymmetry was reported [16] in a confined geometry.
+In the experiment, we developed an original experimental system, in which a metal disk entrains air into viscous liquid in a confined geometry, to find the sheet-forming regime with breakup, in which the entrained air eventually detaches from the disk with forming a bubble: before the detachment a non-axisymmetric constriction region appears in the interface where a sheet of air is formed, which thins down to breakup.
+In the second study, by using the same original experimental system, we further revealed another regime, the corn-forming regime without breakup.
+In this regime, a bubble is not created: the constriction point does not appear in the interface, but a corn of air, which is actually non-axisymmetric, is formed instead of a sheet near the detachment point, as a result of further confining the system [17].
+(We remark here that these regimes were found in the pre-detachment dynamics, while the post-breakup dynamics in the sheet-forming regime has recently been studied [18].) Here, we find a third regime, the corn-forming regime with breakup, in the pre-detachment dynamics, in which the constriction region appears and then a bubble is created but a corn of air, instead of a sheet, is formed at the breaking tip, by investigating different ranges of confining parameters.
+In other words, we find that the transition from breakup to non-breakup can occur within the corn-forming regime, while in any cases the constriction region appears for a breakup but does not for non-breakup: the appearance of the constriction point is a sign of the breakup.
+We further explore, in the third regime, analogy with critical phenomena in thermodynamic transitions [19, 20].
+We show that the master curve, or scaling function for the breakup case is distinctively different from that for the non-breakup case with identifying three scaling exponents and find the exponents for the breakup case dependent on a length scale, which is a continuous number.
+This dependence is in contrast with (I) our previous studies [16, 17] and (II) the standard critical phenomena such as the ferromagnetic transition.
+This is because (I) in our previous studies the corresponding scaling exponents were found to be always the same value even if some length scales are changed and (II) in the standard critical phenomena the exponents usually depend on not continuous but discrete numbers, such as the number of components of the order parameter $n$, representing the symmetry, and dimensionality $d$.
+Since, in critical phenomena, such dependence is known to categorize universality classes, the dependence on a continuous number in the preset hydrodynamic analog indicates the existence of an uncountably infinite number of universality classes, as in some exotic cases known in critical phenomena.
+The deep analogy revealed in the present study suggests a promising avenue for the future study of self-similarity in PDEs by exploring many examples in natural phenomena and industrial processes, in which confined geometries are relevant.
 Figure 1: (a) Experimental setup.
-A metal disk of thickness $D_{0}$ ($=1$ mm) and radius $R$ (101010~{}10to $12.5$ mm) falls in the cell of thickness $D$ ($2$ to $6$ mm) filled with a viscous liquid of kinematic viscosity $\nu$ (100 to 1000 cS).
+A metal disk of thickness $D_{0}$ (=1=1 mm) and radius $R$ (1010~to $12.5$ mm) falls in the cell of thickness $D$ (22 to 66 mm) filled with a viscous liquid of kinematic viscosity $\nu$ (100 to 1000 cS).
 The disk entrains air into the liquid, which finally detaches from the disk.
 The difference between $D$ and $D_{0}$ defines the liquid film thickness $e$.
-(b) Snapshots just before and at breakup illustrating the setting of axes, in the case with topology change for $(R,D_{0},e,\nu)=(10,3,0.5,100)$ in mm or cS.
-Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to detachment of air from the disk for $R=10$ mm, $D_{0}=1$ mm, $\nu=100$ cS.
+(b) Snapshots just before and at breakup illustrating the setting of axes, in the case with topology change for $(R,D_{0},e,\nu)=(10,3,0.5,100)$ in $m$ or cS.
+Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to detachment of air from the disk for $R=10$ mm, $D_{0}=1$ $m$, $\nu=100$ cS.
 The liquid film thickness $e$ are 1, 1.5, and 2 mm, respectively, in (a) to (c).
 The time label 0 ms corresponds to $t=t_{d}$ defined in the text.
-In the three rightmost front-view photos reveal that a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a): ”the critical breakup” lies between $e=$ $1$ and 1.5 mm.
+In the three rightmost front-view photos reveal that a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a): the breakup transition point lies between $e=$ 11 and 1.5 mm.
 Side-view shots near the detachment (0 ms) are separated by 6 ms, but are not synchronized with the front-view shots (a set of side-view snapshots are obtained from an experiment performed on a day different from the day on which the corresponding front-view snapshots but conducted for the same parameters).
@@
 1 (a), we show our experimental setup with explanation.
-The ranges of characteristic lengths, the radius and thickness of disk and the cell thickness, are as follows: $R=10-12.5$ mm, $D_{0}=1$ mm, $D=2-6$ mm.
+The ranges of characteristic lengths, the radius and thickness of disk and the cell thickness, are as follows: R=1012.5R=10-12.5 mm, $D_{0}=1$ $m$, D=26D=2-6 mm.
 The cell width and height are much larger than the length scales $R,D_{0}$, and $D$ (typically 9 and 12 cm, respectively).
@@
 The images are analyzed with Image J and self-made Python codes.
+In the previous studies on the pre-detachment dynamics [16, 17], the sheet-forming with breakup and the corn-forming regime without breakup are typically observed for $D_{0}>\kappa^{-1}$ and $D_{0}<\kappa^{-1}$ at $e=0.5$ mm respectively, where $\kappa^{-1}=\sqrt{\gamma/(\rho g)}\simeq 1.8$ (In the study on the post-breakup dynamics [18], the case of the sheet-forming regime with breakup was studied).
+In the present study on the pre-detachment dynamics, we set $D_{0}$ to a fixed value 1 mm and change $e$ to find the transition from the breakup to the non-breakup in the corn-forming regime.
 II.2 Shape of the air-liquid interface $h(z,t)$
@@
 The region where the width is thin, the interface is almost flat.
-We track the right and left inner edges of the thick curve and describe them by $x=h(z,t)$ and $-h(z,t)$, respectively, except for the case of $e=0.5$ mm, for which we track instead the outer edges.
+We track the right and left inner edges of the thick curve and describe them by $x=h(z,t)$ and h(z,t)-h(z,t), respectively, except for the case of $e=0.5$ mm, for which we track instead the outer edges.
 We determine whether we track the inner or outer edge such that we can have a better collapse of shapes at different time after rescaling, which we discuss below.
 We call the minimum of the function $h(z,t)$ with respect to $z$ ”the constriction point” at which $(x,z)=(h_{m}(t),z_{m}(t))$, i.e., $h_{m}(t)=h(z_{m}(t),t)$.
-II.3 Dynamics sensitive to the film thickness $e$
+II.3 Dynamics dependent on the film thickness $e$
 In Fig.
-2, we show snapshots of air-detachment transitions near the air-solid contact for $e=1.0$, 1.5, and 2 mm, with the other parameters, $R$, $D_{0}$, and $\nu$ fixed (we have tried to analyze the case of $e=2.5$ mm, only to fail because the inner edge was not clear).
+2, we show snapshots of air-detachment near the air-solid contact for $e=1.0$, 1.5, and 2 mm, with the other parameters, $R$, $D_{0}$, and $\nu$ fixed (we have tried to analyze the case of $e=2.5$ mm, only to fail because the inner edge was not clear).
 In three rightmost front-view photos (labeled as 6 ms), detachment without topological change is observed for $e=1$ mm but that with topological change (i.e., breakup) for $e=1.5$ and $2$ mm: a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a).
-In other words, ”the critical breakup” lies between $e=$ $1$ and 1.5 mm.
+In other words, ”the transition point between the breakup and non-breakup” or the breakup transition point lies between $e=$ 11 and 1.5 mm.
 The origin ”0 ms” of the time label ($t-t_{d}$) given in ms in Fig.
@@
 In the previous studies [16] and [17], the cases of $D_{0}\gtrsim\kappa^{-1}$ and $D_{0}\lesssim\kappa^{-1}$ with $\kappa^{-1}=\sqrt{\gamma/(\rho g)}\simeq 1.8$ mm are distinguished by the terms sheet-forming and corn-forming detachments, and the former and latter cases are respectively regarded as the cases with and without topology transition and with and without appearance of the constriction point.
-However, a careful examination on the $e$-dependence in the present study reveals that the previous distinction is applicable only for limited cases: in the present case, we observed both the cases with and without topology transition for $D_{0}\lesssim\kappa^{-1}$ and detachments shown in Fig.
-2 are difficult to categorize by the sheet-forming or corn-forming case because it looks like the corn-forming case but the constriction point does appear.
+However, in the present case, as seen in Fig.
+2, both the cases with and without topology transition are observed for $D_{0}\lesssim\kappa^{-1}$ while the tip form seen from the side-view snapshots looks like a corn rather than a sheet and the constriction point, which is a sign of the breakup, can appear.
+In other words, we find that the transition from breakup to non-breakup can occur within the corn-forming detachment.
 II.4 Dynamics of characteristic length scales
@@
 3 (a1) to (a4), we show the dynamics of three characteristic length scales, $2h_{m}(t)$, $z_{m}(t)$, and $z_{G}(t)$, for different film thickness $e$.
-The quantity $z_{G}(t)$ is the position of the center of gravity of the disk measured from that at the critical time $t=t_{c}$, which slightly differs from $t_{d}$ [$z_{G}(t_{c})=0$ by definition].
+The quantity $z_{G}(t)$ is the position of the center of gravity of the disk measured from that at the critical time $t=t_{c}$, which slightly differs from $t_{d}$ $z_{G}(t_{c})=0$ by definition].
 The critical time $t_{c}$ is determined by extrapolation as the time when $h_{m}$ becomes zero.
@@
 See the text for the determination of the exponent $\delta$.
-$\begin{array}[c]{|c||c||c||c||c|}\hline\cr e\text{ (mm)}&\beta&\Delta&\delta&\Delta/(\beta\delta)\\ \hline\cr\hline\cr 0.5&1.0\pm 0.0&0.97\pm 0.06&1.01\pm 0.01&0.97\pm 0.06\\ \hline\cr 1.0&1.0\pm 0.0&0.79\pm 0.02&0.64\pm 0.00&1.2\pm 0.03\\ \hline\cr 1.5&0.93\pm 0.00&0.78\pm 0.01&0.61\pm 0.00&1.4\pm 0.02\\ \hline\cr 2.0&0.90\pm 0.01&0.77\pm 0.03&0.59\pm 0.00&1.5\pm 0.06\\ \hline\cr\end{array}$
+[…]
 We carefully selected the range of fitting indicated by a pair of cross marks for $h_{m}$ and $z_{m}$ to avoid including the data with large error bars and to use only the data with small error bars, by noting that (i) the error coming from the time resolution tends to become large near $t=t_{c}$ and (ii) the error tends to be large in $z_{m}$ compared with $h_{m}$.
-These points are visualized by the error bars shown for the data at $e=1.0$ mm in (b2) and (b3): The error bars in the horizontal axis are based on the time resolution ±$\pm 0.5$ ms at 2000 fps; The error bars in the vertical axis are on the spacial resolution corresponding to 1 pixel of the images for $h_{m}$ but possibly corresponding to several pixels for $z_{m}$ because of the elongation explained below.
+These issues are visualized by the error bars shown for the data at $e=1.0$ mm in (b2) and (b3): The error bars in the horizontal axis are based on the time resolution $\pm 0.5$ ms at 2000 fps; The error bars in the vertical axis are on the spacial resolution corresponding to 1 pixel of the images for $h_{m}$ but possibly corresponding to several pixels for $z_{m}$ because of the elongation explained below.
 In accordance with Remark (i) and (ii), we can confirm the following in (b2) and (b3): the error bars in the horizontal axis is visible near $t=t_{c}$, while the error bars in the vertical axis are not visible for $h_{m}$ but significant for small $z_{m}$.
@@
 The constriction region tends to elongate near the breakup and the thin thread of air thus formed detaches from the disk when $h_{m}$ is not zero (see Fig.
-6 below).
+6 in Appendix A).
 Because of this elongation of the constriction region the determination of $z_{m}(t)$ tends to become difficult near breakup, which is reflected in larger standard deviation for $\Delta$ (compared with $\beta$) in Tab.
 1.
-As for the fitting of $z_{G}$, we justify the above linear fitting by Fig.
-3 (c), which supports the following scaling law for $z_{G}(t)$ in agreement with the result established in a similar parameter range [11]:
+As for the fitting of $z_{G}$, Fig.
+3 (c) further justifies the linear fitting demonstrated in Fig.
+3 (b1), supporting the following scaling law for $z_{G}(t)$ in agreement with the result established in a similar parameter range [17]:
 $z_{G}(t)=v_{G}t^{\prime}\text{ for }t<t_{c},$ (4)
@@
 Interface shapes after rescaling are reasonably collapsed well on a master curve, especially near the constriction point, as seen in (a2).
-The collapse is significant for shapes before $t-t_{d}=-15$ ms as shown in (a3).
-Figure 4: Shape change for $e=1$, 1.5, 2.0, and 0.5 mm for $(R,D_{0},\nu)=(10,1,100)$ in mm or cS.
+The collapse is significant for shapes before $t-t_{d}=-9$ ms as shown in (a3).
+Figure 4: Shape change for $e=1$, 1.5, 2.0, and 0.5 mm for $(R,D_{0},\nu)=(10,1,100)$ in $m$ or cS.
 Since the right and left interfaces are the mirror image of the other, only the right interface after averaging is shown.
@@
 The time development of interfacial shape before [after] rescaling is respectively shown in (a1) [(a2) and (a3)] for $e=1$ mm.
-We see a better collapse for times not too close to $t=t_{c}$ (before $-9$ ms) [see (a3)].
-The corresponding plots for $e=$1.5, 2.0, and 0.5 mm are shown respectively in (b) to (d).
+We see a better collapse for times not too close to $t=t_{c}$ (before 9-9 ms) [see (a3)].
+The corresponding plots for e=e=1.5, 2.0, and 0.5 mm are shown respectively in (b) to (d).
 The data labeled $R^{\ast}$ in (a3) are obtained for different parameter set $(R,D_{0},\nu)=(12.5,1,100)$, which collapse well on the master curve [the good collapse seems limited for the upper branch (of our focus), i.e., the shape above the constriction point].
@@
 As $e$ increases the constriction region tends to elongate near breakup as announced, which can be confirmed by comparing (b1) with (c1) (see also Fig.
-6 below).
-This elongation is related to the above-mentioned criticality of breakup: topology does change for $e=1.5$ and 2 mm with accompanied by breakup but it does not for $e=1.0$ mm.
+6 in Appendix A).
+This elongation is related to the above-mentioned breakup transition point: topology does change for $e=1.5$ and 2 mm with accompanied by breakup but it does not for $e=1.0$ mm.
 It seems that topological change (i.e., breakup) tends to elongate the constriction region near $t=t_{d}$.
@@
 In Fig.
-4 (a3), even these data obtained for different $R$ collapse well on the master curve, which demonstrates the master curve is independent from (i.e., universal for a change in) $R$ for a fixed $e$, indirectly representing a good reproducibility of the experiment.
+4 (a3), even these data obtained for different $R$ collapse well on the master curve, which demonstrates the master curve is independent from (i.e., universal for change in) $R$ for a fixed $e$, indirectly representing a good reproducibility of the experiment.
 In Fig.
 4 (d1), we show the interfacial shape for $e=0.5$ mm, which is fundamentally different from cases of $e$ larger than 0.5 mm.
-In this case, as mentioned above, we need to track the outer edge instead of inner edge of the dark interface to observe collapse of the renormalized shape as seen in (d2) [and more clearly in (d3)].
+In this case, as mentioned above, we need to track the outer edge instead of inner edge of the dark interface to observe collapse of the renormalized shape seen in (d2) [and more clearly in (d3)].
 In addition, the constriction point does not appear [the point $(h_{m},$ $z_{m})$ is defined as the right contact point (at $y=0$) of the interface with the disk] but the tip forms a corn.
@@
 which scales as $\Gamma(\xi)\sim\xi^{1/\delta}$ for $\xi\gg 1$. This form and the exponent $\delta$ will be quantified below.
+Figure 5: Translated shape functions for $e=1.0,1.5,2.0$, and 0.5 mm on linear scales in the period of good collapse [(a1) to (d1)] and on log-log scales for 21-21 to 30-30 ms with a fitting line obtained by fitting the data in the region between the two black crosses for 21-21 ms [(a2) to (d2)].
+For all cases, $(R,D_{0},\nu)=(10,1,100)$ in $m$ or cS.
 II.6 Identical shapes translating in time
@@
 1, which shows $\delta$ is systematically dependent on $e$ (In the previous study [17], the data corresponding at $e=0.5$ mm are fit by a curve with $\delta=0.75$, which is consistent with the present data: if we fit the data in the region between the pair of blue cross marks, we obtain $\delta=0.75$, which shows the present fitting focus on a region of small errors).
-Figure 5: Translated shape functions for $e=1.0,1.5,2.0$, and 0.5 mm on linear scales in the period of good collapse [(a1) to (d1)] and on log-log scales for $-21$ to $-30$ ms with a fitting line obtained by fitting the data in the region between the two black crosses for $-21$ ms [(a2) to (d2)].
-For all cases, $(R,D_{0},\nu)=(10,1,100)$ in mm or cS.
 The exponent $\delta$ in Eq.
@@
 $(u_{s}(t,z),v_{s}(t,z))\equiv(u(t,x=h(t,z),z),v(t,x=h(t,z),z))$ (12)
-is independent of spacial coordinates. From this, we obtain the spacially constant velocity as the speed of the constriction point:
+is independent of spacial coordinates. This expression means that the spacially constant velocity should be equal to the speed of the constriction point:
 $(u_{s},v_{s})=(\dot{h}_{m},\dot{z}_{m}),$ (13)
@@
 III Discussion
-III.1 Universality and separation of scales
-We show in Fig.
-4 that the shape dynamics in the present parameter range is self-similar in limited ranges of time and space by demonstrating the collapse onto a master curve $\Gamma(\xi)$ by rescaling specified in Eq.
-(6), where $\Gamma(\xi)$ is dependent on $e$.
-From the viewpoint of separation of scales, we expect the present $\Gamma(\xi)$ is independent from $R$, as supported by the data labeled by $R^{\ast}$ in Fig.
-4 (a3).
-This is because among the independent characteristic lengths $e$, $D_{0}$, and $R$, $R$ ($\sim 1$ cm) is the largest scale and thus is well separated from the characteristic scales for the self-similar dynamics.
-We may also expect $\Gamma(\xi)$ is independent from $\eta$ because the master curve $\Gamma(\xi)$ in the previous studies [10, 11] is always independent from $\eta$.
-In fact, we confirmed the master curve obtained for $(e,R,D_{0},\nu)=(0.5,10,1.5,1000)$ well collapses on the master curve in Fig.
-4 (b3) obtained for $(e,R,D_{0},\nu)=(0.5,10,1.0,100)$ near the constriction point (in the region around $z/z_{m}=0$ to 1.4), which suggests the independence of $\Gamma(\xi)$ from $\eta$ and $D_{0}$.
-The independence from $D_{0}$ of $\Gamma(\xi)$ suggested here is also natural from the viewpoint of separation of scales since $D_{0}$ is the second largest scale.
-The independence from $R$ and $\eta$ of the $e$-dependent $\Gamma(\xi)$ suggested here is in accord with the dimensional analysis given below.
-III.2 Comparison with the bubble breakup in non-confined geometry
+III.1 Comparison with the bubble breakup in non-confined geometry
 The translation of the interface in time with keeping the same shape observed in the present study is a property shared by the breakup of a bubble in non-confined geometry as explained below.
@@
 Accordingly, the shape can be described in the cylindrical coordinate $(r,z,\theta)$ by $r=h(t,z)$, where the shape function $h(t,z)$ is independent of the $\theta$ coordinate from the axisymmetry.
-When a high viscosity $\eta$ of the surrounding fluid competes with capillary characterized by the surface tension $\gamma$, the dynamics is known to be described by the following equation (e.g., see Sec.
+When a high viscosity $\eta$ of the surrounding fluid competes with capillarity characterized by the surface tension $\gamma$, the dynamics is known to be described by the following equation (e.g., see Sec.
 9.3.1 of [4]):
@@
 This means that the shape is just translating in time in the horizontal direction as in the present case but with no translation in the vertical direction.
-III.3 Dimensional analysis
+III.2 Dimensional analysis
 We can physically understand how the scaling structure in Eq.
@@
 (1) and (4) at $e=0.5$ mm.
-If we assume Eq.
-(18) is valid even for $e=1.0$ to 2.0 mm, we understood the remaining dependence of the master curve $\Gamma$ on a small length scale $e$ (and possibly on $D_{0}$, which is comparable to $e$) from separation of scales.
-In the time domain in which the collapse is observed in Fig.
-4, while $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ are large enough, $e/z_{m}$ is not large enough so that the $e$-dependence remains in $\Gamma$.
-The present dimensional argument thus suggests the independence from $R$ and $\eta$ of the $e$-dependent $\Gamma(\xi)$ as announced above.
-However, to justify Eq.
-(18) even for $e=1.0$ to 2.0 mm, we need to know the parameter dependence of the coefficients $c_{1}$ and $c_{2}$ in Eq.
-(2), which we will explore in a future study.
-III.4 Analogy with critical phenomena
+In this way, we can naturally understand the self-similar structure at $e=0.5$ mm based on a dimensional argument.
+In the case of $e=1.0$ to 2.0 mm, in order to justify a similar argument, different from the case of $e=0.5$ mm, we have yet to know the parameter dependence of the coefficients $c_{1}$ and $c_{2}$ in Eq.
+(2).
+Such a study would be interesting especially in terms of universality and separation of scales.
+For example, if Eq.
+(18) is still valid even for $e=1.0$ to 2.0 mm, we may understand universality of $\Gamma$ for change in $R$ but not in $e$ indicated in the present experiment: In the time domain in which the collapse is observed in Fig.
+4, while $R/z_{m}$ and $l/z_{m}\simeq 1/t^{\prime 2}$ are large enough compared with the length scale over which the collapse is observed, $e/z_{m}$ is not large enough so that the $e$-dependence remains in $\Gamma$.
+In other words, the dependence on length scales could disappear if the scale is well separated from the length scale of collapse, but could remain if not.
+This issue of universality and separation of scales is explored for the post-breakup dynamics [18].
+Similar study for the present pre-detachment dynamics would be an interesting future problem, although this requires careful experiments.
+Note that experiments on the pre-detachment is technically more challenging because the pre-detachment dynamics tends to be more sensitively affected by the entry conditions.
+We remark that the above tentative argument for $e=1.0$ to 2.0 mm suggests universality of $\Gamma$ for change in $\eta$.
+In fact, the master curve in the previous studies on the pre-detachment dynamics [16, 17] is always independent of $\eta$.
+III.3 Analogy with critical phenomena
 In critical phenomena in thermodynamic transitions, various thermodynamic quantities exhibit power laws near the critical point.
@@
 Furthermore, the critical exponents satisfy scaling relations such as $\alpha+2\beta+\gamma=2$ and such relations are universally satisfied by the critical exponents of any universality classes.
-Such remarkable features can be understood based on the scaling hypothesis as explained in Appendix, and the scaling hypothesis naturally emerges in the renormalization group theory [12, 13].
+Such remarkable features can be understood based on the scaling hypothesis as explained in Appendix B, and the scaling hypothesis naturally emerges in the renormalization group theory [19, 20].
 The scaling hypothesis in the magnetic case can be expressed as
@@
 (2).
-Time $t$ ($=t-t_{c}$) and position $z$ ($=z-z_{c}$) correspond to temperature $T$ and external field $H$, respectively, for ”the order parameter $h(t,z)$,” a quantity which is non-zero before $t_{c}$ but is zero after $t_{c}$ at $z=0$: the pre-breakup dynamics of our focus corresponds to the temperature range below $T_{c}$.
+Time $t$ ($t-t_{c}$) and position $z$ ($z-z_{c}$) correspond to temperature $T$ and external field $H$, respectively, for ”the order parameter h(t,z)h(t,z),” a quantity which is non-zero before $t_{c}$ but is zero after $t_{c}$ at $z=0$: the pre-detachment dynamics of our focus corresponds to the temperature range below $T_{c}$.
 The master curve $\Gamma(\xi)$ corresponds to the function $\Psi(\xi)$, which is called the scaling function and defines the critical exponent $\delta$ through $\Psi(\xi)\simeq\xi^{1/\delta}$ for $\xi\gg 1$ (and thus $\delta$ in Eq.
@@
 (19), where the exponents $\beta$, $\Delta$, and $\delta$ in critical phenomena satisfy Eq.
-(10) as shown in Appendix.
+(10) as shown in Appendix B.
 In other words, the critical exponents $\beta$, $\Delta$, and $\delta$ in the present case are the ones summarized in Tab.
@@
 (10) is more violated as $e$ increases, and that that the exponents are monotonic functions of the thickness $e$.
-In addition, we may expect from the viewpoint of separation of scales and the dimensional analysis discussed above the exponents are universally shared by the systems with different $R$ and $\eta$ for a given $e$ (at $D_{0}=1$ mm).
+In addition, we may expect from the above arguments on the universality of $\Gamma(\xi)$, which defines one of the exponents $\delta$, the exponents are universally shared by the systems with different $R$ (and $\eta$) for a given $e$ (at $D_{0}=1$ mm).
 Namely, the present universality class may be determined by a continuous parameter $e$.
 This is in contrast with typical critical phenomena, in which universality class is determined by discrete parameters such as $n$ and $d$, although examples of critical exponents depending on continuous parameters are known when marginal operators play a role, which includes non-linear diffusion equations [22].
-In the present case, a renormalization group analysis suggests a different scenario, which will be discussed elsewhere.
 IV Conclusion
-We investigate the pre-breakup dynamics of a bubble surrounded by a viscous liquid in a confined geometry focusing on the dependence on the lubricating film thickness $e$ and find that the dynamics is self-similar and that the scaling exponents characterizing the self-similarity is sensitively dependent on $e$.
-We elucidate similarity between the present singular dynamics and critical phenomena, showing that the counterpart of the critical exponents are dependent on a continuous variable $e$ and could be shared by a universality class formed by the systems with different $R$ and $\eta$ for a fixed $e$.
+We investigate the pre-detachment dynamics of a bubble surrounded by a viscous liquid in a confined geometry focusing on the dependence on the lubricating film thickness $e$ for a fixed $D_{0}$ (=1.0=1.0 mm).
+The present results are in contrast with our previous studies on the pre-detachment dynamics [16, 17] in that (I) we add a third regime, the corn-forming regime with breakup, to the previously found two regimes, the sheet-forming regime with breakup and the corn-forming regime without breakup, and that (II) we find the scaling exponents characterizing the self-similarity are dependent on a length.
+We elucidate similarity between the present singular dynamics and critical phenomena, showing that the counterpart of the critical exponents are dependent on a continuous variable $e$ and could be shared by a universality class formed by the systems with different $R$ (and $\eta$) for a fixed $e$ (at $D_{0}=1.0$ mm).
+The dependence on the continuous variable suggest the existence of an uncountably infinite number of universality classes.
+The physical origin of the dependence on $e$ but not on $R$ could be understood from separation of scales, as explored in the study of the post-breakup [18].
 The analogy between the singular dynamics and critical phenomena elucidated in the present study underscores the importance of exploring symmetry in the future study of the singular dynamics by exploiting confined geometries for the following reasons.
@@
 (2) The singular dynamics is widely observed in nature and a variety of confined geometries are increasingly important in industrial applications and natural phenomena, from microfluidics [27, 28], petroleum industry [29], to geology [30].
-In fact, the analogy with critical phenomena can be deepened further by developing a renormalization group analysis for the present singular dynamics, which will be discussed elsewhere.
+In fact, it is recently shown that the analogy with critical phenomena can be deepened further by developing a renormalization group analysis for the breakup of a bubble in the non-confined geometry [31], which will be further explored elsewhere for the present case.
+Figure 6: Snapshots obtained at 2000 fps near the detachment point for $e=0.5$ mm (top) and $e=1.0$ mm (bottom) at $t-t_{d}=-1.5$ to +1.5+1.5 ms, respectively.
+The snapshots at the center correspond to $t-t_{d}=-0$ ms, where ”the order parameter” $h_{m}$ is not precisely zero.
 Acknowledgements.
 This work was supported by JSPS KAKENHI Grant Number JP19H01859 and JP24K00596.
-Appendix A Scaling relations in critical phenomena
-Figure 6: Snapshots obtained at 2000 fps near the detachment point for $e=0.5$ mm (top) and $e=1.0$ mm (bottom) at $t-t_{d}=-1.5$ to $+1.5$ ms, respectively.
-The snapshots at the center correspond to $t-t_{d}=-0$ ms, where ”the order parameter” $h_{m}$ is not precisely zero.
+Appendix A Dynamics close to the breakup time
+We compare in Fig.
+6 the dynamics close to $t=t_{c}$, which shows qualitative differences for $e=0.5$ and 1.0 mm.
+In the former, sharp tip is formed, which detaches from the disk.
+In the latter, the constriction region elongates near $t=t_{c}$, forming a thin thread of air, which detaches from the disk when $h_{m}$ is not zero.
+Appendix B Scaling relations in critical phenomena
 We summarize here the basic knowledge on critical phenomena, focusing on the temperature range below $T_{c}$ ($T<T_{c}$).
@@
 1.
-Appendix B Dynamics close to the breakup time
-We compare in Fig.
-6 the dynamics close to $t=t_{c}$, which shows qualitative differences for $e=0.5$ and 1.0 mm.
-In the former, sharp tip is formed, which detaches from the disk.
-In the latter, the constriction region elongates near $t=t_{c}$, forming a thin thread of air, which detaches from the disk when $h_{m}$ is not zero.